Let f and g be continuous functions from a metric space X to metric space Y such that they agree on a dense subset of a X .then prove that f=g


1
Expert's answer
2021-07-27T14:45:06-0400

Let "M=\{x\in X: f(x)=g(x)\}\subset X". Since f and g are continuous functions from a metric space X to metric space Y, M is a closed subset (see below), i.e. "\overline{M}=M". By condition, M is dense, i.e. "\overline{M}=X". Therefore, "M=X" and "f=g" everywhere. Q.E.D.


Let's show that M is a closed subset. Let "x_n\in M" be any sequence in M, converging to "x\in X".

That is, "x=\lim\limits_{n\to\infty}x_n" . Since f and g are continuous, then "f(x)=\lim\limits_{n\to\infty}f(x_n)=\lim\limits_{n\to\infty}g(x_n)=g(x)". This means that "x\in M" and that M is closed.